OpenStudy (anonymous):

Simplify: 18 – 4 + 12 ÷ 2

OpenStudy (anonymous):

13

OpenStudy (anonymous):

how did you not get 8?

OpenStudy (anonymous):

I got 13..

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

You subtract 18-4 so its 14...then add 12 you get 26 then divide it by 2...then you get 13..its really simple.

OpenStudy (anonymous):

that doesnt make sense if you use PEMDAS or BOMDAS??!!

ganeshie8 (ganeshie8):

you should get 20

OpenStudy (anonymous):

By BODMAS , 18-4+12/2 .. would be : 18 - 4 + 6 = 24 - 4 = 20 .. No ??

OpenStudy (anonymous):

thats what i thought but i did it on a quiz on online school and got it wrong. im confused!

OpenStudy (anonymous):

Now i'm confused too :/